OpenStudy (pottersheep):

Is (+-i√15)/5 equal to +-i√(3/5) ???? Shouldnt i be over 5???

OpenStudy (turingtest):

\[\frac{\sqrt{15}}5=\frac{\sqrt{3\cdot5}}5=\sqrt{\frac35}\]

OpenStudy (pottersheep):

Thanks, shouldn'tthe i still be over 5 tho?

OpenStudy (turingtest):

nope...

OpenStudy (turingtest):

\[\pm i\frac{\sqrt{15}}5=\pm i\frac{\sqrt{3\cdot5}}5=\pm i\sqrt{\frac{3\cdot5}{25}}=\pm i\sqrt{\frac35}\]

OpenStudy (pottersheep):

oo but the i was over 5 to begin with

OpenStudy (turingtest):

but the 5 can move around...

OpenStudy (pottersheep):

ok let me try to process this xD

OpenStudy (turingtest):

\[\pm\frac i5\sqrt{15}=\pm i\frac{\sqrt{15}}5=\pm i\frac{\sqrt{3\cdot5}}5=\pm i\sqrt{\frac{3\cdot5}{25}}=\pm i\sqrt{\frac35}\]

OpenStudy (turingtest):

remember that\[\frac{2i}3=i\frac23\]not\[\frac{2i}3=\frac i3\frac23\]

OpenStudy (turingtest):

\[a\frac bc=b\frac ac=\frac{ab}c\]

OpenStudy (pottersheep):

:O it is???

OpenStudy (pottersheep):

ooo

OpenStudy (turingtest):

what else could it be?

OpenStudy (pottersheep):

THANKSSS

OpenStudy (turingtest):

\[\frac{ab}c=a\cdot\frac bc\neq\frac ac\cdot\frac bc\]

OpenStudy (turingtest):

welcome :D